Enter the value for x that makes the equation –2x – 1.5(4 – x) = 20 – 3x true.


Answer by Boreal(15235)

–2x – 1.5(4 – x) = 20 – 3x
-2x-6+1.5x=20-3x
-.5x-6=20-3x
2.5x=26
x=10.4
----------------
-20.8-1.5(-6.4)=20-31.2
-20.8+9.6=-11.2, and this is correct